OpenStudy (anonymous):

Solve Ln(x-5) +ln(x+2)=ln(x-14)

OpenStudy (tkhunny):

Have you considered logarithm functions? \(log(a) + log(b) = log(ab)\)

OpenStudy (anonymous):

I am lost :(

OpenStudy (tkhunny):

You should find yourself. This is the simplest logarithm transformation. If you cannot do this one,. you ARE lost. Sorry. Just the truth. This is the rule: \(\log(a) + log(b) = log(ab)\) This turns addition of logarithms into the logarithm of a product. Look REALLY HARD at your problem statement. Do you see ANYTHING like that?